OpenStudy (anonymous):

When simplifying: 32*50 is the answer: 4/10 or 4/100 (all numbers are in brackets)

OpenStudy (anonymous):

\[\sqrt{32}*\sqrt{50} = 4\sqrt{10} or 4\sqrt{100}\]

OpenStudy (anonymous):

What does 32 x 50 equal?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

1600 so the 4 squared is the sixteen and times 10 squared is 1600 or times 100

OpenStudy (anonymous):

40*40=1600

OpenStudy (anonymous):

Does 4 times the square root of 10 = 1600? Or does 4 times the square root of 100 equal 1600?

OpenStudy (anonymous):

the square root of 100 can be simplified further

OpenStudy (anonymous):

4^2 *10^2=1600